OpenStudy (anonymous):

There is a population of Drosophila flies in a small fruit shop. Due to sudden mutation, the eye color of some of the flies changes from red to white. Red (represented by R) is the dominant gene, and white (represented by r) is a recessive gene. Due to this, the gene frequency of allele r changes from 0.38 to 0.40. What will be the new genotype frequencies in that population? ( p + q = 1, p2 + 2pq + q2 = 1) A. p2 = 0.36, 2pq = 0.48, q2 = 0.16 B. p2 = 0.32, 2pq = 0.23, q2 = 0.45 C. p2 = 0.25, 2pq = 0.36, q2 = 0.39

OpenStudy (aaronq):

if r=q=0.4 then R=p=1-0.4=0.6 just plug those in into p^2, 2pq, and q^2

OpenStudy (anonymous):

Could you explain that alittle bit to me?

OpenStudy (aaronq):

Well just differentiate, gene frequencies versus population frequencies. when they say gene frequencies, they mean q and r , and you know that p+q=1 so if they give you q or p, then q=1-p, right? similarly if they give you population frequencies (phenotypes), they mean p^2+2pq+q^2=1 p^2 is for the homozygous dominant phenotypes (RR) 2pq is for the heterozygotes phenotypes (Rr) q^2 is for the homozygous recessive phenotypes (rr)

OpenStudy (tesa12345):

@oOKawaiiOo i need help with the correct answer on this please

OpenStudy (ookawaiioo):

give me 4 mins

OpenStudy (tesa12345):

okay thanks

OpenStudy (ookawaiioo):

answers is A

OpenStudy (ookawaiioo):

but let me explain....give me a min

OpenStudy (tesa12345):

thank you!!!(:

OpenStudy (ookawaiioo):

p + q = 1 I desingated p as red (dominant trait= R) and q as white (recessive= r) Since the r (recessive) changed from 0.38 to 0.40. p + q = 1 ; remember q is r (recessive) p + 0.4 = 1 ; so p (dominant = R) is 0.6 p^2 + 2pq + q^2 = 1 (0.6)^2 + 2pq + (0.40)^2 = 1 (0.36) + 2pq + (0.16) = 1 ; so 2pq is 0.48 Answer A) p2 = 0.36, 2pq = 0.48, q2 = 0.16

OpenStudy (ookawaiioo):

desinated* not desingated LOL

OpenStudy (ookawaiioo):

designated* FML